A boy is 42 years younger than his father. In 8 years time, he will be 1/4 times as old as his father .Finde their present age.
Answer by scott8148(6628)

"A boy is 42 years younger than his father" ___ b = f - 42

"In 8 years time, he will be 1/4 times as old as his father" ___ b + 8 = (1/4)(f + 8) ___ 4b = f - 24

subtract the equations to eliminate f